ON THE HOMOLOGY OF CONFIGURATION SPACES C((M, M_{o}) × R^{n}; X)

Jie Wu

Abstract. The homology with coefficients in a field of the configuration spaces C(M × R^{n}, M_{o} × R^{n}; X) is determined in this paper.

Typeset by AMS-TEX